OpenStudy (anonymous):

0.5,2.5,4.5,6.5.... Medals wbg

OpenStudy (anonymous):

@thefoe1990

OpenStudy (anonymous):

In addition of 2 0.5 + 2= 2.5 + 2= 4.5 + 2 =6.5

OpenStudy (anonymous):

your o a role @kandylovesmath